Can you please help me write out an equation for the word problem?
If a plane can travel 400 miles per hour into the wind and 540 miles with the wind, find the speed of the wind and the speed of the plane in still air.

Answer by scott8148(6628)

"400 miles per hour into the wind" ___ p - w = 400

"540 miles with the wind" ___ p + w = 540

adding the equations ___ 2p = 940 ___ p = 470

substituting ___ 470 + w = 540 ___ w = 70